### v1

@@ Abstract
Computational Scientific Discovery; Synthetic Biology; Active Learning; Inductive Logic Programming; Matrix; Relational Algebra.
@@ 1 Introduction
A priority in scientific research is the development of autonomous systems to enhance the scientific discovery process. Currently, most endeavours in this domain concentrate on either laboratory automation or machine learning applied to scientific datasets. Sub-symbolic approaches usually demand a large number of annotated data and extensive trial-and-error. It is often difficult to modify the structure of these machine learning systems or extract human-interpretable information. The application of these systems in an experimental workflow is limited by data efficiency, flexibility and interpretability. There is a resurgence of interest in incorporating AI in experimental automation and scientific understanding simultaneously. In this work, we investigate an integration of techniques for planning experiments, optimising resource utilisation and interpretable machine learning to support biological discovery.
@@
Our study explores the applicability of the abductive reasoning and active learning approach to comprehensive biological systems. Given that biological relationships are commonly described logically, ILP is particularly adept at operating on biological knowledge bases. The integration of abductive reasoning and active learning via ILP was successfully demonstrated in the context of biological discovery by the prominent Robot Scientist King04:RobotScientist . The Robot Scientist performed active learning by strategically selecting key experiments to achieve more data- and cost-effective gene function learning than random experimentation. However, this demonstration was limited to only 17 genes in the aromatic amino acid pathway of yeast. In the present study, we advance this approach by looking at Escherichia coli (E. coli). The iML1515 genome-scale metabolic model (GEM) iML1515 encompasses 1515 genes and 2719 metabolic reactions of the E. coli strain K-12 MG1655, a versatile host organism for metabolic engineering to produce specific compounds. Notably, the iML1515 model represents a 90-fold increase in complexity compared to the aromatic amino acid pathway investigated by the Robot Scientist.
@@
Despite the extensive mapping of the E. coli genome to metabolic functions in iML1515, it was shown to yield inaccurate phenotype predictions bernstein_evaluating_2023 . A classical method to address errors in metabolic pathways involves auxotrophic growth experiments, where specific genes are deleted to render the organism incapable of synthesizing essential compounds beadle_genetic_1941 . For accurate simulation of auxotrophic mutant experiments, the computational representation of GEMs must faithfully depict the underlying network structure and facilitate efficient exploration of metabolic pathways.
@@
Considering these requirements, Petri nets emerge as suitable modelling tools for GEMs. Petri nets are a class of directed bipartite graphs with diverse applications in Computer Science and Biology reisig_understanding_2013 ; sahu_advances_2021 . Petri nets can represent a structured knowledge base while capturing dynamic behaviours in a system. In contrast, logic programs allow reasoning over relationships between entities but cannot directly reflect the changes in concepts and entity states. By integrating logic programs with Petri nets, we enable a more holistic approach to learning biological systems that combines dynamic simulations with logical reasoning.
@@
Figure 1: The hypotheses and instances are represented as logic programs to describe gene-enzyme associations and auxotrophic mutant experiment conditions respectively. $BMLP_{active}$ encodes the GEM iML1515 and simulates the auxotrophic mutant experiment outcomes. It actively consults a data source to request ground truth labels that minimise the expected experimental cost when pruning hypotheses. The underlying BMLP iteratively refutes gene function hypotheses inconsistent with labelled training examples. While $BMLP_{active}$ to date only learns from synthetic data, a laboratory robot can be integrated to perform selected experiments.
@@
Learning complex biological systems like GEMs presents significant challenges in terms of computational complexity and empirical feasibility. To address these challenges, we present a new system $BMLP_{active}$ (Figure 1) based on a novel approach called Boolean Matrix Logic Programming (BMLP) and an implementation in SWI-Prolog. We use logic programs to encode the biochemical, genetic, and genomic relationships in iML1515. Logical inferences to simulate phenotypic effects are accelerated by boolean matrices via BMLP. This speeds up the runtime of phenotypic effect simulations by a factor of 170 compared to SWI-Prolog without BMLP on a single computation thread. We implement active learning to select auxotrophic mutant experiments that minimise the expected experimental resource cost. Empirically, we delete gene functions in iML1515 and use $BMLP_{active}$ to learn hypotheses to recover the deleted gene-enzyme associations. $BMLP_{active}$ successfully restores these removed gene functions through active learning while reducing 90% of the experimental resource cost required by randomly selected experiments. It also lowers the number of experimental data needed to recover these gene functions compared to random experiment selection, reducing the sample complexity. When operating within a finite budget, our approach can deliver optimal experimental outcomes, whereas studies employing random experiment selection might not reach completion.
@@
Furthermore, we apply $BMLP_{active}$ to the challenge of learning digenic functions, which holds significant implications for drug development and therapeutic interventions costanzo_global_2016 . These digenic interactions are related to isoenzymes - two enzymes responsible for the same reaction. The intricate relationship between complex genotypes and phenotypes remains largely unexplored in most organisms costanzo_global_2019 . Additionally, some digenic interactions are dynamic and can vary depending on the growth conditions bernstein_evaluating_2023 . A comprehensive understanding of digenic interactions would expedite strain engineering efforts. We show that $BMLP_{active}$ can converge to the correct gene-isoenzyme function with as little as 20 training examples. This shows a significant promise for $BMLP_{active}$ to address complex genetic interactions on the whole genome level.
@@
    We describe the setting of Boolean Matrix Logic Programming (BMLP).
@@
    We provide a SWI-Prolog implementation of a BMLP algorithm that improves the runtime of phenotypic effect simulations in a genome-scale metabolic network model (GEM) by 170-fold compared to simulations written only in SWI-Prolog.
@@
    We built a BMLP system called $BMLP_{active}$ to actively select experiment instances for learning gene-enzyme associations in a GEM.
@@
    We present a lower bound of the sample complexity ratio between active learning and random experiment sampling with empirical evidence to support it.
@@
    We empirically show that $BMLP_{active}$ reduces 90% experimental resource cost and the number of experiments needed for learning gene-enzyme associations compared to randomly sampled experiments.
@@
    We empirically demonstrate that $BMLP_{active}$ can be devised to learn gene-isoenzyme associations in GEM with a faster convergence to the correct gene-isoenzyme function than randomly sampled experiments.
The rest of this paper is organised as follows. In Section 2, we summarise work on machine learning of genome-scale metabolic networks, matrix algebra, Petri nets and active learning. In Section 3, we introduce key concepts in matrix algebra and an elementary class of Petri nets that we use to model a GEM, called one-bounded elementary nets (OENs). In Section 4, we describe a transformation from an OEN to a subclass of recursive datalog programs. In Section 5, we introduce our BMLP framework to evaluate this class of datalog. In Section 6, we describe our active learning approach based on prior work ase_progol ; hocquette_how_2018 ; progol . In Section 7, we explain the $BMLP_{active}$ implementation for encoding iML1515, GEM simulation and active learning gene functions. Section 8 includes our empirical results on simulation runtime, phenotype prediction, active learning gene functions and isoenzymes. In Section 9, we explain the current limitations of $BMLP_{active}$ and discuss future work.
@@ 2.1 Learning genome-scale metabolic network models
2.2 Matrix algebraic approaches to logic programming.
@@
Obtaining the least Herbrand model of recursive datalog programs can be reduced to computing the transitive closure of boolean matrices peirce_collected_1932 ; copilowish_matrix_1948 . Fischer and Meyer fischer_boolean_1971 study a logarithmic divide-and-conquer computation technique by viewing relational databases as graphs and show a significant computational improvement over the “naive” iteration of matrix operators. A similar approach is explored in ioannidis_computation_1986 for computing the fixpoint of recursive Horn clauses. An ILP system called DeepLog muggleton_hypothesizing_2023 employs boolean matrices for learning recursive datalog programs. It repeatedly computes square boolean matrices to quickly choose optimal background knowledge for deriving target hypotheses. Propositional logic programs can be represented as boolean matrices to describe the dependencies in gene regulatory networks as boolean networks sato_boolean_2021 ; sato_differentiable_2023 . The boolean networks can be learned by minimizing a differentiable cost function. However, none of these techniques is designed for Petri nets simulation of GEMs. In our framework, we encode a GEM as a Petri net and use boolean matrices to evaluate a large datalog program that simulates the Petri net.
@@
Recent work primarily studies the mapping of logic programs to linear equations over tensor spaces. SAT problems are investigated in lin_satisability_2013 under a linear algebra framework. Truth values of domain entities, logical relations and operators for first-order logic can be evaluated using tensors-based calculus grefenstette_towards_2013 . Computing algebraic solutions can approximate recursive relations in first-order Datalog sato_linear_2017 . Based on this approach, abduction can be performed by encoding a subset of recursive datalog programs in tensor space sato_abducing_2018 . Tensor representations allow differentiable infrastructures such as neural networks to perform probabilistic logical inferences cohen_tensorlog_2020 . This can be done using dynamic programming to propagate beliefs for stochastic first-order logic programs. The probabilistic parameters can be optimised using deep learning frameworks yang_differentiable_2017 . Limitations have been identified for solving certain recursive programs due to the difficulty of computing some arithmetic solutions whereas these can be solved by iterative bottom-up evaluations sato_linear_2017 . In addition, algebraic approximations are usually calculated in linear algebraic methods where the correctness of the solutions cannot be guaranteed. Our approach does not have these issues since it uses boolean matrices for iterative bottom-up computation.
2.3 Petri nets in logic programming.
@@
Little attention has been paid to representing and simulating Petri nets with logic programs. With logic programming, techniques to execute, analyse and learn Petri nets can be implemented in a single programming language domenici_petri_1990 . The early work domenici_petri_1990 describes places and transitions of a Petri net in terms of n-ary ground facts. Srinivasan et al. srinivasan_knowledge_2012 attempt to learn n-ary ground facts that represent Petri net transitions in the ILP ILP1991 system Aleph aleph . Few studies have looked at encoding Petri nets by Answer Set Programming (ASP) gelfond_stable_1988 . Behrens and Dix behrens_model_2007 use Petri nets for analysing interactions in multi-agent systems with ASP programs. Anwar et al. anwar_encoding_2013 provide an ASP encoding of Petri nets with coloured tokens to model the behaviours of biological systems. This encoding is adapted by Dimopoulos et al. dimopoulos_encoding_2020 for simulating reversible Petri nets that allow transitions to fire in both directions. These systems all use high-level symbolic representations of Petri nets as deductive knowledge bases. Logic programming approaches rely on navigating background knowledge to determine the consistency of hypotheses. However, the runtime of deductive inference in large-scale knowledge bases limits the applicability of these systems.
2.4 Active learning in computational scientific discovery
Active learning techniques provide a machine learning algorithm with more control over the training examples it uses cohn_improving_1994 . A typical approach to active learning is via membership queries where the machine learner queries a data source to label instances according to some strategy. An active learner can be implemented to explore the version space which contains all hypotheses that are expressible by a hypothesis language and are consistent with a set of training examples mitchell_generalization_1982 . Instances are chosen according to a measure of informativeness such as entropy shannon_mathematical_1963 , diameter tosh_diameter_based_2017 , the size mitchell_generalization_1982 ; dasgupta_coarse_2005 and the shrinkage hocquette_how_2018 of the version space for eliminating competing hypotheses. For a binary classification task, when learners are allowed to look at $n$ unlabelled instances to compare, an active learner has $n$ times larger probability of selecting an instance with maximal entropy than a passive learner hocquette_how_2018 . The general principle for a binary active learner is to label instances closest to being consistent with half of the hypotheses in the version space mitchell_generalization_1982 . These instances reject up to half of the hypothesis space regardless of their classifications. In the optimal case when all instances are maximally discriminative, the target hypothesis can be found via a binary search with depth logarithmic of the hypothesis space size angluin_queries_2001 . $BMLP_{active}$ makes membership queries but our objective is to not only reduce the number of labels but also minimise the expected cost of experimentation.
Computational systems such as the BACON systems langley_scientific_1987 , the Lagramge systems todorovski_integrating_2006 , recent equation discovery systems brunton_discovering_2016 ; guimera_bayesian_2020 ; petersen_deep_2020 are primarily designed to formulate symbolic hypotheses from experimental results. However, one cannot directly employ these systems for experimental planning. The automation of experimental design was investigated by the pioneering Robot Scientist King04:RobotScientist as an endeavour to bring together logical reasoning, active learning and laboratory automation. The Robot Scientist automatically proposes explanatory hypotheses, actively devises validation experiments, performs experiments using laboratory robotics, and interprets empirical observations. Abduction and active learning are supported by ASE-Progol ase_progol based on the ILP system Progol progol . Experiments were actively selected to minimise the expected cost of experimentation for learning gene functions in the aromatic amino acid synthesis pathway of yeast. It was combined with a laboratory robot to automatically reconstruct the yeast metabolic pathway model and the active learning strategy significantly outperformed random experiment selection King04:RobotScientist . Later, a high-throughput laboratory automation platform, Adam king_automation_2009 extended this approach by constructing a logical model of yeast consisting of 1200 genes. In comparison, we apply $BMLP_{active}$ to a genome-scale metabolic network, which is significantly larger than the models in King04:RobotScientist ; king_automation_2009 .

### v3

@@ Abstract
Computational Scientific Discovery; Synthetic Biology; Active Learning; Inductive Logic Programming; Matrix.
@@ 1 Introduction
A priority in scientific research is the development of autonomous systems to enhance the scientific discovery process. Currently, most efforts in this domain focus on either laboratory automation or machine learning applied to scientific datasets. Sub-symbolic approaches usually demand a large number of annotated data and extensive trial-and-error. It is often difficult to modify the structure of these machine learning systems or extract human-interpretable information. The application of these systems in an experimental workflow is limited by data efficiency, flexibility and interpretability. There is a resurgence of interest in incorporating AI in experimental automation and scientific understanding simultaneously. In this work, we investigate an integration of techniques for planning experiments, optimising resource utilisation and interpretable machine learning to support biological discovery.
@@
Our study explores the applicability of abductive reasoning and active learning to comprehensive biological systems. Given that biological relationships are commonly described logically, ILP is particularly adept at operating on biological knowledge bases. The integration of abductive reasoning and active learning via ILP was successfully demonstrated in the context of biological discovery by the prominent Robot Scientist King04:RobotScientist . The Robot Scientist performed active learning by strategically selecting key experiments to achieve more data- and cost-effective gene function learning than random experimentation. However, this demonstration was limited to only 17 genes in the aromatic amino acid pathway of yeast. In the present study, we advance the scope of logic-based gene function learning by looking at Escherichia coli (E. coli). The genome-scale metabolic model (GEM), iML1515 iML1515 , encompasses 1515 genes and 2719 metabolic reactions of the E. coli strain K-12 MG1655, a versatile host organism for metabolic engineering to produce specific compounds. Notably, the iML1515 model represents a 90-fold increase in complexity compared to the aromatic amino acid pathway investigated by the Robot Scientist.
@@
Despite the extensive mapping of the E. coli genome to metabolic functions in iML1515, it was shown to yield inaccurate phenotype predictions bernstein_evaluating_2023 . A classical method to address errors in metabolic pathways involves auxotrophic growth experiments, where specific genes are deleted to render the organism incapable of synthesizing essential compounds beadle_genetic_1941 . For accurate simulation of auxotrophic mutant experiments, the computational representation of GEMs must faithfully depict the underlying network structure and effectively facilitate the exploration of metabolic pathways.
@@
Considering these requirements, Petri nets emerge as suitable modelling tools for GEMs. Petri nets are a class of directed bipartite graphs with diverse applications in Computer Science and Biology reisig_understanding_2013 ; sahu_advances_2021 . Petri nets can represent a structured knowledge base while capturing dynamic behaviours in a system. In contrast, logic programs allow reasoning over relationships between entities but cannot directly reflect the changes in concepts and entity states. To integrate logic reasoning with Petri nets, we explore a more holistic approach to learning metabolic networks that combines recursive logic programs with pathway reachability analysis.
@@
Figure 1: The hypotheses and instances are represented as logic programs to describe gene functions (gene-enzyme associations) and auxotrophic mutant experiment conditions respectively. $BMLP_{active}$ encodes the GEM iML1515 and simulates the auxotrophic mutant experiment outcomes. It actively consults a data source to request ground truth labels that minimise the expected experimental cost based on a user-provided cost function when pruning hypotheses. The underlying BMLP iteratively refutes gene function hypotheses inconsistent with labelled training examples. While $BMLP_{active}$ to date only learns from synthetic data, a laboratory robot can be integrated to perform selected experiments.
@@
Learning complex biological systems like GEMs presents significant challenges in terms of computational complexity and empirical feasibility. To address these challenges, we present a new system $BMLP_{active}$ (Figure 1) based on a novel approach called Boolean Matrix Logic Programming (BMLP) and an implementation in SWI-Prolog. We use logic programs to encode the biochemical, genetic, and genomic relationships in iML1515. While simulating auxotrophic mutant experiments, logical inferences to predict phenotypic effects are accelerated by boolean matrices via BMLP. This speeds up the runtime for predicting phenotypic effects by a factor of 170 compared to SWI-Prolog without BMLP on a single computation thread. We implement active learning to select auxotrophic mutant experiments that minimise the expected experimental cost via a user-defined cost function. Empirically, we delete gene functions in iML1515 and use $BMLP_{active}$ to learn hypotheses to recover the deleted gene functions. $BMLP_{active}$ successfully restores these removed gene functions through active learning while reducing 90% the optional nutrient substance cost required by randomly selected experiments. It additionally lowers the number of experimental data needed to recover these gene functions compared to random experiment selection, reducing the sample complexity. When operating within a finite budget, our approach can deliver optimal experimental outcomes, whereas studies employing random experiment selection might not reach completion.
@@
Furthermore, we apply $BMLP_{active}$ to learning digenic functions, which holds significant implications for drug development and therapeutic interventions costanzo_global_2016 . These digenic interactions are related to isoenzymes - two enzymes responsible for the same reaction. The intricate relationship between complex genotypes and phenotypes remains largely unexplored in most organisms costanzo_global_2019 . Additionally, some digenic interactions are dynamic and can vary depending on the growth conditions bernstein_evaluating_2023 . A comprehensive understanding of digenic interactions would expedite strain engineering efforts. We show that $BMLP_{active}$ can converge to the correct gene-isoenzyme mapping with as little as 20 training examples. This shows a significant promise for $BMLP_{active}$ to address complex genetic interactions on the whole genome level.
@@
    We describe a novel problem of Boolean Matrix Logic Programming (BMLP).
@@
    We implement a BMLP algorithm, BMLP-IE, in SWI-Prolog to predict phenotypic effects by computing reachability queries in GEMs based on a subclass of Petri net, which empirically improves simulation runtime by 170-fold compared to SWI-Prolog without BMLP-IE.
@@
    We built a BMLP system, $BMLP_{active}$, which uses BMLP-IE to actively select experiment instances for learning gene functions in a GEM.
@@
    We present a theoretical sample complexity ratio between active learning and random experiment sampling and empirically highlight the effect of actively selecting examples in reducing the total training examples needed.
@@
    We empirically show that $BMLP_{active}$ reduces 90% optional nutrient substance cost and the number of experiments needed for learning gene functions compared to randomly sampled experiments.
@@
    Our empirical results demonstrate that $BMLP_{active}$ can be devised to learn gene-isoenzyme associations in GEM with a faster convergence to the correct gene-isoenzyme mapping than randomly sampled experiments.
The rest of this paper is organised as follows. In Section 2, we summarise work on machine learning of genome-scale metabolic networks, matrix algebra, Petri nets and active learning. In Section 3, we introduce key concepts in matrix algebra and an elementary class of Petri nets that we use to model a GEM, called one-bounded elementary nets (OENs). In Section 4, we describe a transformation from an OEN to a subclass of recursive logic programs, linear and immediately recursive datalog. In Section 5, we introduce our BMLP framework to evaluate this class of logic programs. In Section 6, we describe our active learning approach based on prior work ase_progol ; progol . In Section 7, we explain the $BMLP_{active}$ implementation for encoding iML1515, GEM simulation and active learning gene functions. Section 8 includes our empirical results on simulation runtime, phenotype prediction, active learning gene functions and gene-isoenzyme mapping. In Section 9, we explain the current limitations of $BMLP_{active}$ and discuss future work.
@@ 2 Related work
2.1 Active learning in computational scientific discovery
Active learning techniques provide machine learning algorithms with more control over the training examples they use cohn_improving_1994 . A typical approach to active learning is via membership queries where the machine learner queries a data source to label instances according to some strategy. An active learner can be implemented to explore the version space which contains all hypotheses that are expressible by a hypothesis language and are consistent with a set of training examples mitchell_generalization_1982 . Instances are chosen according to a measure of informativeness such as entropy shannon_mathematical_1963 , diameter tosh_diameter_based_2017 , the size mitchell_generalization_1982 ; dasgupta_coarse_2005 and the shrinkage hocquette_how_2018 of the version space for eliminating competing hypotheses. For a binary classification task, when learners are allowed to look at $n$ unlabelled instances to compare, an active learner has $n$ times larger probability of selecting an instance with maximal entropy than a passive learner hocquette_how_2018 . The general principle for a binary active learner is to label instances closest to being consistent with half of the hypotheses in the version space mitchell_generalization_1982 . These instances reject up to half of the hypothesis space regardless of their classifications. In the optimal case when all instances are maximally discriminative, the target hypothesis can be found via a binary search with depth logarithmic of the hypothesis space size angluin_queries_2001 . $BMLP_{active}$ makes membership queries but our objective is to not only reduce the number of labels but also minimise the expected cost of experimentation.
Computational systems such as the BACON systems langley_scientific_1987 , the LAGRANGE systems todorovski_integrating_2006 , recent equation discovery systems brunton_discovering_2016 ; guimera_bayesian_2020 ; petersen_deep_2020 are primarily designed to formulate symbolic hypotheses from experimental results. However, one cannot directly employ these systems for experimental planning. The automation of experimental design was investigated by the pioneering Robot Scientist King04:RobotScientist as an endeavour to bring together logical reasoning, active learning and laboratory automation. The Robot Scientist automatically proposes explanatory hypotheses, actively devises validation experiments, performs experiments using laboratory robotics, and interprets empirical observations. Abduction and active learning are supported by ASE-Progol ase_progol based on the ILP system Progol progol . Experiments were actively selected to minimise the expected cost of experimentation for learning gene functions in the aromatic amino acid synthesis pathway of yeast. It was combined with a laboratory robot to automatically reconstruct the yeast metabolic pathway model and the active learning strategy significantly outperformed random experiment selection King04:RobotScientist . Later, a high-throughput laboratory automation platform, Adam king_automation_2009 extended this approach by constructing a logical model of yeast consisting of 1200 genes. In comparison, we apply $BMLP_{active}$ to a genome-scale metabolic network, which is significantly larger than the models in King04:RobotScientist ; king_automation_2009 .
@@ 2.2 Learning genome-scale metabolic network models
2.3 Matrix algebraic approaches to logic programming
@@
In our BMLP framework, we use boolean matrices to evaluate fixpoints of large recursive datalog programs. Obtaining the least Herbrand model of recursive datalog programs can be reduced to computing the transitive closure of boolean matrices peirce_collected_1932 ; copilowish_matrix_1948 . Fischer and Meyer fischer_boolean_1971 study a divide-and-conquer boolean matrix computation technique by viewing relational databases as graphs. A similar approach is explored by Ioannidis ioannidis_computation_1986 for computing the fixpoint of recursive Horn clauses. An ILP system called DeepLog muggleton_hypothesizing_2023 employs boolean matrices for learning recursive datalog programs. It repeatedly computes square boolean matrices to quickly choose optimal background knowledge for deriving target hypotheses. The mapping from biological knowledge bases to propositional logic programs is explored in sato_boolean_2021 ; sato_differentiable_2023 where logic programs are represented as boolean matrices by considering gene regulatory networks as boolean networks. In contrast, we encode a GEM as a Petri net and a recursive datalog program for computing reachability queries.
@@
Recent work primarily studies the mapping of logic programs to linear equations over tensor spaces. SAT problems are investigated in lin_satisability_2013 under a linear algebra framework. Truth values of domain entities, logical relations and operators for first-order logic can be evaluated using tensors-based calculus grefenstette_towards_2013 . Computing algebraic solutions can approximate recursive relations in first-order Datalog sato_linear_2017 . Based on this approach, abduction can be performed by encoding a subset of recursive datalog programs in tensor space sato_abducing_2018 . Tensor representations allow differentiable infrastructures such as neural networks to perform probabilistic logical inferences cohen_tensorlog_2020 . This can be done using dynamic programming to propagate beliefs for stochastic first-order logic programs. The probabilistic parameters can be optimised using deep learning frameworks yang_differentiable_2017 . Limitations have been identified for solving certain recursive programs due to the difficulty of computing some arithmetic solutions whereas these can be solved by iterative bottom-up evaluations sato_linear_2017 . In addition, algebraic approximations are usually calculated in linear algebraic methods where the correctness of the solutions cannot be guaranteed. However, our BMLP approach does not have these issues since it uses boolean matrices for iterative bottom-up computation.
2.4 Petri nets in logic programming
@@
Little attention has been paid to representing and simulating Petri nets with logic programs. With logic programming, techniques to execute, analyse and learn Petri nets can be implemented in a single programming language domenici_petri_1990 . The early work domenici_petri_1990 describes places and transitions of a Petri net in terms of n-ary ground facts. Srinivasan et al. srinivasan_knowledge_2012 attempt to learn n-ary ground facts that represent Petri net transitions in the ILP ILP1991 system Aleph aleph . Few studies have looked at encoding Petri nets by Answer Set Programming (ASP) gelfond_stable_1988 . Behrens and Dix behrens_model_2007 use Petri nets for analysing interactions in multi-agent systems with ASP programs. Anwar et al. anwar_encoding_2013 provide an ASP encoding of Petri nets with coloured tokens to model the behaviours of biological systems. This encoding is adapted by Dimopoulos et al. dimopoulos_encoding_2020 for simulating reversible Petri nets that allow transitions to fire in both directions. These systems all use high-level symbolic representations of Petri nets as deductive knowledge bases. Logic programming approaches rely on navigating background knowledge to determine the consistency of hypotheses. However, non-high-throughput deductive inferences limit these systems’ applicability for dealing with increasingly complex knowledge bases such as GEMs. $BMLP_{active}$ addresses this limitation by utilising a boolean matrix encoding of GEMs and boolean matrix operations to accelerate inferences.